# The fate of poor man’s Majoranas in the long Kitaev chain limit

### v1

@@ Abstract
Conductance signatures: Zero-bias conductance is often used as a signature to search for PMMs in a minimal Kitaev chain Dvir et al. (2023); Zatelli et al. (2023); Ten Haaf et al. (2024); van Driel et al. (2024); Bordin et al. (2024a). We use the Python package Kwant Groth et al. (2014) to calculate the conductance $G_{\alpha\beta}=dI_{\alpha}/dV_{\beta}$, attaching leads to the two outer QDs and introducing the coupling $t_{l}$ between the QDs and the leads. Varying $\mu_{L}$ and $\mu_{R}$, for different values of $\mu_{M}$, we expect a clear signature in the ROT associated with PMMs. Within the ROT, there is a crossing of the conductance at $\mu_{L}=\mu_{R}$. As $\mu_{M}$ is varied away from its value in the ROT, the zero-energy conductance exhibits an anti-crossing and the non-local conductance changes its sign Tsintzis et al. (2022); Dvir et al. (2023); Zatelli et al. (2023); Ten Haaf et al. (2024). We find this signature for both parameter sets, see Fig. 4. This means that zero-energy conductance cannot clearly distinguish false PMMs from true PMMs, i.e., those that evolve into topological states in the long chain limit.
@@
Figure 5: Finite-energy conductance of true PMMs versus false PMMs. Normalized conductance at energy $E$, as the chemical potential of the left QD, $\mu_{L}$, is varied, while the chemical potential of the right QD remains at its value in the ROT. (a) and (c) show the local conductance $G_{LL}$, (b) and (d) show the nonlocal conductance $G_{LR}$. For each setup all conductance values are normalized with respect to the maximum value. The parameters for (a) and (b) are the same as for (a)-(c) in Fig. 2, with $t_{l}/\Delta=0.01$. The parameters for (c) and (d) are the same as for (d)-(f) in Fig. 2, with $t_{l}/\Delta=0.006$. Both sets of parameters result in qualitatively similar results, but only (a) and (b) are related to true MBSs in the long chain limit, whereas (c) and (d) are related to trivial states in a long chain.
@@
Another signature which has been used to identify PMMs is the finite energy conductance when tuning the chemical potential of only one, e.g. the left, QD away from the ROT Dvir et al. (2023); Ten Haaf et al. (2024); Bozkurt et al. (2024). However, we again find that there is no clear qualitative difference in this signature between a false and a true PMM, see Fig. 5. In the SM sm , we also show finite energy conductance results when both QDs are tuned away from the ROT. This, too, is inadequate to distinguish false from true PMMs.
@@
Origin of false PMMs: As shown in Fig. 3, in the long chain limit, the false PMMs discussed here are related to degenerate energy levels that are split from the bulk states. These false PMMs can be associated with localized zero-energy states in a system consisting of only two QDs, see SM sm , in particular, these false PMMs arise because the first and last normal QDs of a chain couple directly to only one superconducting QD, but the bulk normal QDs couple directly to two superconducting QDs.
@@
Conclusion: We found that signatures so far associated with PMMs can also originate from “false PMMs”, which are states that are near-zero-energy and highly localized in the minimal chain, but do not evolve into topological states in the long chain limit. Although we focussed on minimal chains here, we find also that short uniform chains consisting of a few QDs Bordin et al. (2024a) can also host false PMMs, but with a reduced excitation gap, see SM sm . It is therefore interesting to consider if true and false PMMs can be distinguished without having to construct a long uniform chain. Finally, throughout we assumed uniform parameters in the long chain limit, however, it is important to emphasise that this is a theoretical tool to determine the nature of PMMs. In realistic experimental chains there is no reason to assume uniform parameters. For instance, although gates can easily tune chemical potential, it is not expected that other parameters such as $g$-factor, SOI strength and direction, onsite $U$, and many more can be uniform, this could cause problems with the identification of MBSs in the longer chains Hess et al. (2022, 2023b); Legg et al. (2024).

### v2

@@ Abstract
Conductance signatures: Zero-bias conductance is often used as a signature to search for PMMs in a minimal Kitaev chain Dvir et al. (2023); Zatelli et al. (2023); Ten Haaf et al. (2024); van Driel et al. (2024); Bordin et al. (2024a). We use the Python package Kwant Groth et al. (2014) to calculate the conductance $G_{\alpha\beta}=dI_{\alpha}/dV_{\beta}$, attaching leads to the two outer QDs and introducing the coupling $t_{l}$ between the QDs and the leads. Varying $\mu_{L}$ and $\mu_{R}$, for different values of $\mu_{M}$, we expect a clear signature in the TR associated with PMMs. Within the TR, there is a crossing of the conductance at $\mu_{L}=\mu_{R}$. As $\mu_{M}$ is varied away from its value in the TR, the zero-energy conductance exhibits an anti-crossing and the non-local conductance changes its sign Tsintzis et al. (2022); Dvir et al. (2023); Zatelli et al. (2023); Ten Haaf et al. (2024). We find this signature for both parameter sets, see Fig. 4. Therefore, zero-energy conductance cannot clearly distinguish true from false PMMs.
@@
Figure 5: Finite-energy conductance of true PMMs versus false PMMs. Normalized conductance at energy $E$, as the chemical potential of the left QD, $\mu_{L}$, is varied, while the chemical potential of the right QD remains at its value in the TR. (a) and (c) show the local conductance $G_{LL}$, (b) and (d) show the nonlocal conductance $G_{LR}$. For each setup all conductance values are normalized with respect to the maximum value. The parameters for (a) and (b) are the same as for (a)-(c) in Fig. 2, with $t_{l}/\Delta=0.01$. The parameters for (c) and (d) are the same as for (d)-(f) in Fig. 2, with $t_{l}/\Delta=0.006$. Both sets of parameters result in qualitatively similar results, but only (a) and (b) are related to true MBSs in the long chain limit, whereas (c) and (d) are related to trivial states in a long chain.
@@
Another signature which has been used to identify PMMs is the finite energy conductance when tuning the chemical potential of only one, e.g. the left, QD away from the TR Dvir et al. (2023); Ten Haaf et al. (2024); Bozkurt et al. (2024). However, we again find that there is no clear qualitative difference in this signature between a false and a true PMM, see Fig. 5. In the SM sm , we also show finite energy conductance results when both QDs are tuned away from the TR. This, too, is inadequate to distinguish false from true PMMs.
@@
Origin of false PMMs: As shown in Fig. 3, in the long chain limit, the false PMMs discussed here are related to degenerate energy levels that are split from the bulk states. These false PMMs can be associated with localized zero-energy states in a system consisting of only two QDs, see SM sm . In particular, these false PMMs arise because the first and last normal QDs of a chain couple directly to only one superconducting QD, but the bulk normal QDs couple directly to two superconducting QDs.
@@
False PMMs in longer chains: Although we focused here on minimal chains with just two QDs, we also find that chains consisting of a few QDs Bordin et al. (2024a); ten Haaf et al. (2024) can host Majorana-like states – which we call “scaled PMMs” – that do not evolve into MBSs in the long chain limit, see SM sm . This demonstrates that the generalisation of the PMM concept to finite chains also does not guarantee that these states will eventually evolve into MBSs in the long chain limit, even in setups where the chain can be uniformly extended.
Conclusion: We have demonstrated that the observation of PMMs in minimal Kitaev chains is neither necessary nor sufficient to ensure that these states evolve into MBSs in longer (uniform) chains and conductance signatures so far associated with such states can also originate from “false PMMs”, i.e., highly localized near-zero-energy states in the minimal chain that evolve into trivial states in the long chain limit. One possibility to distinguish true from false PMMs might be the study of excited states in a longer chain, see SM sm . However, throughout we assumed uniform parameters in the long chain limit and it is important to emphasise that this is a theoretical tool to determine the nature of PMMs. In realistic experimental chains, there is no reason to assume uniform parameters. For instance, although gates can tune chemical potentials, it is not expected that other parameters such as $g$-factor, SOI strength and direction, onsite $U$, and many more can be uniform, which could cause problems with the identification of MBSs in longer chains Hess et al. (2022, 2023b); Legg et al. (2024). It is therefore interesting to consider if true and false PMMs can be distinguished before a long uniform chain is assembled.